# Neurodevelopmental disorders modeling using isogeometric analysis, dynamic domain expansion and local refinement

### v1

@@ 6 Experimental human iPSC-derived and rat hippocampal neuron cultures
In this section, we outline the protocols and details for culturing human induced pluripotent stem cells (iPSCs) and rat hippocampal neuron cultures. Conducting experimental cultures to obtain neuron growth observations is necessary for computational modeling as it provides a reference to simulations and for comparison with realistic results. This approach allows us to better understand the intricacy of neurite outgrowth, improving the ability of the NDD model to reflect real-world biological behaviors in NDD studies. Human iPSC-derived neurons are selected to better correlate between the NDDs model and actual human neuron growth. Building upon our previous studies of healthy rat hippocampal neurons [32, 68], we now extend our focus to unhealthy rat hippocampal neurons to deepen our understanding of NDDs. In future work, we intend to include unhealthy human iPSC-derived neurons currently being developed at the Mayo Clinic to broaden our investigation of underlying disorder mechanisms.
@@
Rat hippocampal neuron cultures. For apoptotic rat hippocampal neuron culture results used in this paper, are drawn from failed cultures during protocol optimization for our previously published work creating a dataset and model of rat hippocampal neuron growth [32, 33] with slight modification. The neurons, taken from embryonic day 18 and cryopreserved (A36513, Gibco, USA), were grown in dishes coated with poly-D-lysine (P6407, Sigma-Aldrich, USA) following the manufacturer’s guidelines [97]. The seeding density was 10,000 cells per square centimeter in Neurobasal Plus medium (A3582901, Gibco, USA) supplemented with 2% B-27 Plus (A3582801, Gibco, USA). Cultures were maintained at 37 degrees Celsius with 5% CO2. It is common when establishing new neural cultures to optimize media formulation, media change frequency, and the fraction of media replaced during media changes (ADD REF). During this optimization process several cultures underwent apoptosis with neurite beading. Imaging was carried out using an Echo Revolve Microscope (Echo Revolve |||| R4, inverted, BICO, USA) with a 12-megapixel color camera at 20X and 40X magnifications and at intervals of 0.5, 1, 2, 3, 4, and 6 days post-inoculation. The images of experimental rat hippocampal neuron cultures undergoing persumed excitotoxic apoptosis are selected for qualitative comparison with neuron growth simulations affected by NDD parameters.
7 Numerical results and validation with expeirments
@@
In this section, we first explain the implementation specifics of the proposed IGA phase field NDDs model. Then, we showcase simulation results of healthy and unhealthy neurons with several parameter studies designed to investigate NDDs. Finally, we compare our simulation outcomes with experimental data, focusing on cases with external guided-cue mechanisms. This comparative analysis aims to validate the effectiveness and accuracy of our NDDs model in simulating biomimetic neurite behaviors and disorders.
@@
We implemented the IGA NDDs model in C++, leveraging the extensive capabilities and scalabilities of the PETSc library [61]. PETSc as the computational backbone offers a significant computational efficiency advantage over MATLAB-based implementations [31]. This is mainly because the PETSc library is written in compile language and tailored for high-performance computing environments, while MATLAB is a scripting language that faces limitations due to its inherently interpreted mode of execution, making it less efficient for tasks requiring intensive computation. Through the utilization of the MPI for parallel processing, our model facilitates efficient distribution of computations across multiple threads [62, 98], thereby cutting down execution times significantly when compared with MATLAB analysis. Although MPI leverages the immense computational power enabled by supercomputers, communication time among processors becomes a bottleneck. Mesh partitioning using METIS [64] ensures optimal load balancing for the computationally intensive simulations. This is crucial for the scalability and efficiency of the model, particularly when simulating complex neurite structures on locally refined truncated T-splines with a large number of DOFs because mesh alterations during execution often lead to unbalanced loads. Unbalanced loads across threads could lead to unnecessary slowdown and waiting during inter-threads/node communications. We ran simulations on the Bridges-2 supercomputer at Pittsburgh Supercomputing Center [99, 100] using 128-thread regular memory nodes.
@@
For clarity, we provide a detailed list of variables for NDDs and healthy neurons in Table 1. The domain $\phi$ is initialized with a central filled circle with radius $r_{0}$, representing the cell. The initial values of $\theta$ are randomized to between $[0,1]$ across the domain with $T$ set to 0 at the beginning. Due to the significant variability in biophysics and growth behaviors across different types of neurons, the parameter values used in the phase field NDDs model were sourced from established literature [29, 30, 101] and fine-tuned empirically to capture realistic neuron growth. These parameters can be adjusted to better simulate biomimetic growth of specific neuron types  [32]. With these parameter settings, we first simulate healthy neuron growth and analyze single- and multiple-neuron growth scenarios (Section 7.1). Then, we conduct a simulation of abnormal neurite morphological transformation using the NDDs model with three case studies that each focuses on a specific parameter (Section 7.2). Finally, we compare simulation results with experimental observations to validate our NDDs model (Section 7.3).
@@ 7.1 Healthy neuron growth simulation
In our results, the neuron, represented by $\phi$, is shown in yellow, and the surrounding medium is shown in blue. For single neuron cases, the results (Fig. 4A) show that the model captures a broad spectrum of neurite outgrowth behaviors, from axon elongations to complex branching. Concurrently, our multiple-neuron simulations investigate interactions within complex neurite network formation, varying from 2-7 neurons (Fig. 4B). Starting from 2-neuron cases, we gradually increase complexity to configurations with up to 7 randomly placed neurons. These multi-neuron simulations showcase individual neurite growth patterns and the intricate neurite interactions with responses of other neurons in proximity. With the detailed dynamics of $c_{tubu}$ and $c_{neur}$, these biomimetic single and multi-neuron simulation results showcase the potential of our model for in-depth neurodevelopmental research and provide a solid foundation for exploring NDDs.
@@ 7.2 Neurodevelopmental disorders
    Optimal neurotrophin concentration $c_{opti}$. This parameter models the inverse relationship between neurotrophin concentration level and neurite survival and is analyzed for its impact mostly on neurite retraction;
@@
For each study, the parameter being analyzed was adjusted based on the specified ranges listed in Table 1. All other parameters were set according to their default values listed in the same table. By adjusting these parameters in our NDD simulations, we can not only enhance our understanding of the pathophysiological mechanisms of NDDs but also examine the capability of our model to capture their intricate dynamics, which is critical for developing targeted interventions and advancing neurodevelopmental research.
@@
Optimal neurotrophin concentration $c_{opti}$. By adjusting the value of $c_{opti}$ while keeping other parameters consistent with Table 1, we can better understand its impact on neurite morphological transformations in NDDs. We first analyze neurite growth behaviors under different $c_{opti}$ values ranging from $0.5$ to $3.0$ at around $40,000$ iterations (Fig. 5A). As $c_{opti}$ increases from $0.5$ to $1.5$, there is a noticeable improvement in neurite density and complexity, promoting more extended growth patterns. This process indicates a positive response of neurites to optimal neurotrophin levels, matching our previous expectations of enhanced growth behaviors. As $c_{opti}$ progresses from $1.5$ to $3.0$, there is a noticeable shift towards highly branched and extensively multidirectional neurite structures. This range of growth patterns, from minimal to excessive branching, shows a critical threshold level of neurotrophins necessary for neurodevelopmental processes. Notably, lower values of $c_{opti}$ lead to suppressed neurite outgrowth, validating that there is an inverse relationship between $c_{neur}$ and neurite survival [87], as well as a critical balance of $c_{neur}$ for maintaining healthy neurite development. We then explore the dynamic effects of varying $c_{opti}$ in our NDDs model to understand its impact on neurodevelopmental processes. For simulation cases in Fig. 5B, we initialize $c_{opti}$ as 1 to simulate healthy neurite outgrowth and then reduce $c_{opti}$ to 00 at $150,000$ iterations across cases involving different numbers of neurons. The first two rows in Fig. 5B showcase the growth behavior of individual neurons, showing a straightforward process of neurite retraction as $c_{opti}$ decreases, highlighting the sensitivity of single neurons to $c_{neur}$ and $c_{opti}$ levels. The third and fourth rows extend this analysis to multi-neuron configurations with two and three neurons to investigate neurite retraction and interactions among multiple neurons. These simulation results highlight the intricate influence of $c_{opti}$ levels on the morphology transformation of neurites and the potential implications to neurite retractions and atrophy, vital for developing effective therapeutic strategies targeting NDDs.
@@
Neurotrophin diffusion rate $D_{c}$. In this study, we focus on the neurotrophin diffusion rate $D_{c}$ that is responsible for controlling the diffusion of $c_{neur}$ to support $\phi$ interface balance. We analyze the effect of $D_{c}$ with varying magnitude from $1.0$ to $6.0$ on single- and multiple-neuron cases, and results at around $100,000$ iterations that best exhibit atrophy are shown in Fig. 6. The results show severe atrophy when the $D_{c}$ value is below $2.0$. As the $D_{c}$ value approaches $3.0$, atrophy becomes moderate. This atrophy behavior indicates that lower values of $D_{c}$, particularly at 1.0, severely restrict the diffusion necessary for generating a neurotrophin gradient field for Eqn. 13 that stabilizes the phase field interface during neurite outgrowth. This insufficient gradient compromises neurite integrity and leads to the degeneration or atrophy of neurite structures due to inadequate neurotrophic support. As a result, as neurite tips continue growing out, there is not enough concentration to support and maintain the grown neurite structures, leading to neurite atrophy. On the other hand, as $D_{c}$ increases, particularly beyond $5.0$, a noticeable improvement in neurite continuity is observed in both single- (Fig. 6A) and multiple-neuron cases (Fig. 6B), markedly reducing the incidence of neurite atrophy. This highlights the necessity of maintaining a diffusion rate $D_{c}$ to ensure a consistent concentration distribution surrounding the growing neurite, thereby supporting the uninterrupted neurodevelopmental process and potentially mitigating disorder progression in NDDs.
@@
These detailed parameter studies utilizing our NDDs model demonstrate NDDs morphological transformations to specific simulation parameters. The impact of $c_{opti}$ is most dominantly on neurite retraction and branching, the influence of $D_{c}$ leads to neurite atrophy, and the degradation rates $k_{p75}$ and $k_{2}$ affect the thickness of neurite growth. By categorically investigating these parameters, we have identified crucial factors that predominantly influence the progression of NDDs in our computational model. Our analysis provides crucial insights into how variations in these key parameters can trigger or exacerbate the symptoms associated with NDDs. This knowledge is vital in guiding ongoing research toward understanding the pathophysiological mechanisms underlying these disorders. In the long term, this model could enable the development of more targeted therapeutic planning, focusing on the most influential factors of neuron development and treatment strategy. This approach will enhance our understanding of potential NDDs triggering mechanisms and broaden the potential reach for effective treatment strategies by incorporating the effect of parameters with biophysical processes in simulating realistic neurodevelopmental progress.
7.3 Validation with experiements
@@
A crucial component of our NDDs model study involves comparing it with experimental observations. We apply the external-cue guided mechanism [32] to simulate the existence of external attractive cues in the extracellular matrix for several simulation cases and set the results against experimental neuron growth patterns (Fig. 8-10). By selectively placing external cues around the neuron, our model can effectively guide neurite outgrowth toward these external cues, allowing it to capture the complex and dynamic growth patterns observed in experimental cultures. These simulations explore how well our model reflects real-world biological behaviors, thereby validating its effectiveness for studying NDDs. This study compares simulation results with single neuron cultures to better understand individual neurite morphology. Although culturing denser neuron cultures is feasible, the densely overlapping neurite structures will significantly complicate neurite identifications and comparisons. Considering these limitations, efforts have been made to develop sparser neuron cultures to improve the clarity and reliability of neurite analysis.
@@
Figure 8: Comparison of simulated neuron growth patterns and experimental cultures of human iPSC-derived neurons. (A) Comparisons of single neuron cases up to day 5. (B) Comparisons of multiple-neuron cases involving two and three neurons. External cue placements are marked with cyan starts. These comparisons provide a visual validation of NDDs model to replicate key aspects of neurite morphology and developmental processes in real biological environments.
@@
Comparison with healthy human iPSC-derived neurons. First, we validate the model by comparing single-neuron growth patterns from day 1 to day 5 (Fig. 8A). Experimental images of human iPSC-derived neurons cultured over five days are used as the reference. We incorporate the external cue-guided mechanism by placing external cues around the soma to direct neurite outgrowth. These results indicate that the NDDs model can capture the observed neurite growth patterns when external cues are placed according to experimental culture images. Next, we test the NDDs model with multiple neuron configurations (Fig. 8B). This setup involved comparing experimental images of multi-neuron growth with simulations where external cues were selectively placed to guide neurite outgrowth. The simulation outcomes closely aligned with the experimental images, demonstrating that the model can biomimetically capture complex growth behaviors in multi-neuron environments. Although measuring the exact biological conditions and translating them into parameters in the phase field model present inherent challenges, this comparison showcases the effectiveness and potential of our NDDs model in capturing essential aspects of the neurodevelopmental process, highlighting its potential as a powerful tool for studying NDDs. These results validate the biomimetic capabilities and applicability of the NDDs model in simulating realistic neurodevelopmental processes, providing valuable insights for targeted therapeutic strategies.
@@
Comparison with rat hippocampus neurons undergoing excitotoxic apoptosis. Utilizing the same external cue placement strategy, we perform a detailed comparative analysis of two neuron growth cases with disorders by modifying $D_{c}$ and $c_{opti}$ and compared with rat hippocampal neuron cultures undergoing excitotoxic apoptosis [32, 68] (Fig. 9). The NDDs model captures atrophy and retraction by adjusting $D_{c}$, which controls the diffusion of neurotrophin concentration, and $c_{opti}$, which sets the optimal neurotrophin concentration for neurodevelopmental processes. Based on experimental images of neuron growth undergoing excitotoxic apoptosis (Fig. 9A&E). Red outlines are empirically traced based on both experimental observations and accumulated expertise. External cues are placed to guide healthy neurite patterns (Fig. 9B&F). Subsequently, we decrease the magnitude of $D_{c}$ from $6$ to $1$ to simulate reduced diffusion in neurotrophin concentration (Fig. 9C&G). This simulation reveals that neurites suffer significant atrophy and disconnections, showcasing a stark contrast to simulated healthy growth. In the other comparison, we lowered $c_{opti}$ magnitude from $1$ to 00 (Fig. 9D&H). The results show that neurites not only experience significant atrophy but also demonstrate retraction behaviors, which are influenced by the cascading energy balance at the interface. This study highlights the critical influence of $D_{c}$ and $c_{opti}$ on pathological neuron growth affected by NDDs. These parameter adjustments show great potential to enable the NDDs model to simulate a variety of morphological abnormalities and provide insights into the dynamics of NDDs, aiding in the development of future targeted therapies.
@@
Figure 9: Comparison of experimental neuron cultures undergoing excitotoxic apoptosis and simulation results. (A&E) Experimental images of neurons exhibiting disorders are highlighted in red outlines. (B&F) Simulated healthy neurons using the NDDs model. (C&G) Simulated neuron morphology with disorders induced by reducing the diffusion rate $D_{c}$ from 6 to 1. (D&H) Simulated neuron morphology with disorders induced by lowering the optimal concentration $c_{opti}$ from 1 to 0.
Figure 10: Additional comparison of experimental rat hippocampal neurons undergoing excitotoxic apoptosis and simulated growth patterns. (A&B) Showcases single neuron scenarios, illustrating the progression and diversity of growth patterns.
@@
In addition to the above two in-depth comparison studies, we apply our NDDs model to simulate a range of cases following the same approach (Fig. 10). The first and third rows in Fig. 10 are the experimental images of neuron growth undergoing excitotoxic apoptosis (in red outlines). The second and fourth rows in Fig. 10 are the corresponding NDDs model simulation results of NDDs neurite growth patterns. Each simulation is able to simulate biomimetic growth behaviors under disorder conditions with atrophy and retraction, with the influence of external cues guided mechanism on neurites. These results not only demonstrate the capability of our NDDs model but also showcase its effectiveness in simulating realistic NDDs-associated neurite growth observed in experimental cultures.

### v2

@@ 6 Experimental human iPSC-derived and rat hippocampal neuron cultures
In this section, we outline the protocols and details for culturing human induced pluripotent stem cells (iPSCs) and rat hippocampal neuron cultures. Conducting experimental cultures to obtain neuron growth observations is necessary for computational modeling as it provides a reference to simulations and for comparison with realistic results. This approach allows us to better understand the intricacy of neurite outgrowth, improving the ability of the NDDs model to reflect real-world biological behaviors in NDDs studies. Human iPSC-derived neurons are selected to better correlate between the NDDs model and actual human neuron growth. Building upon our previous studies of healthy rat hippocampal neurons [32, 68], we now extend our focus to unhealthy rat hippocampal neurons to deepen our understanding of NDDs. In future work, we intend to include unhealthy human iPSC-derived neurons currently being developed at the Mayo Clinic to broaden our investigation of underlying disorder mechanisms.
@@
Rat hippocampal neuron cultures. For rat hippocampal neuron culture results used in this paper, images are drawn from failed cultures during protocol optimization for our previously published work creating a dataset and model of rat hippocampal neuron growth [32, 33] and from cultures in which cells experienced mechanical damage.
@@
For culture condition optimization examples, cryopreserved embryonic day 18 neurons (A36513, Gibco, USA) were thawed and grown in dishes coated with poly-D-lysine (PDL) (P6407, Sigma-Aldrich, USA) following the manufacturer’s guidelines [97] with slight modification to optimize conditions for low density culture. Primary neurons are notoriously challenging to culture [98]. It is not uncommon when establishing new neural cultures to optimize media formulation, media change frequency, and the fraction of media replaced during media changes. During this optimization process several cultures underwent cell death with neurite breaking and fragmentation. The seeding density for all samples was either 20,000 or 100,000 cells per square centimeter in Neurobasal Plus medium (A3582901, Gibco, USA) supplemented with 2% B-27 Plus (A3582801, Gibco, USA). Cultures were maintained at 37C with 5% $CO_{2}$. Imaging was carried out using an Echo Revolve Microscope (Echo Revolve |||| R4, inverted, BICO, USA) with a 12-megapixel color camera at 20X and 40X magnifications. The images included here for model comparison were all taken 3-4 days after plating.
For samples in which neurons experienced mechanical damage, wells were coated first with PNIPAM following manufacturer protocols and subsequently coated with a PDL-ECM gel mixture (PECM) composed of Geltrex ECM Gel (Thermo Fisher, A1413301) diluted 1:100 in DMEM and mixed with 10 $\mu$g/mL PDL [99, 100] for 1 hour at room temperature. The coating solution was then removed and the coated dishes were left uncovered in the laminar biosaftey cabinet for 2 hours. Cryopreserved embryonic day 18 neurons (A36513, Gibco, USA) were then seeded at a density of 20,000 cells/cm2 and cultured for 2 days at 37C in 5% $CO_{2}$. After the initial culture period cells were lifted from the plate using a gelatin plunger transfer technique [101]. Briefly, a gelatin gel plunger was gently placed in contact with the cells and the cells were released from the plate by incubation at 20C. The gelatin plunger was then removed and placed into a separate PDL coated well. This new plate was briefly incubated at 20C followed by incubation at 37C to liquefy the gelatin. All wells were gently washed with Hank’s balanced salt solution to remove gelatin and imaged.
It should be noted that these conditions were not intended to replicate any specific developmental disorder. They are included here for qualitative comparison to demonstrate the models ability to create degenerating and damaged morphologies.
7 Numerical results and validation with experiments
@@
In this section, we first explain the implementation specifics of the proposed IGA phase field NDDs model. Then, we showcase simulation results of healthy and unhealthy neurons with several parameter studies designed to investigate NDDs. Finally, we compare our simulation outcomes with experimental data, focusing on cases with external cue-guided mechanisms. This comparative analysis aims to evaluate the effectiveness and accuracy of our NDDs model in simulating biomimetic neurite behaviors and disorders.
@@
We implemented the IGA NDDs model in C++, leveraging the extensive capabilities and scalabilities of the PETSc library [61]. PETSc as the computational backbone offers a significant computational efficiency advantage over MATLAB-based implementations [31]. This is mainly because the PETSc library is written in compiled language and tailored for high-performance computing environments, while MATLAB is a scripting language that faces limitations due to its inherently interpreted mode of execution, making it less efficient for tasks requiring intensive computation. Through the utilization of the MPI for parallel processing, our model facilitates efficient distribution of computations across multiple threads [62, 102], thereby cutting down execution times significantly when compared with MATLAB analysis. Although MPI leverages the immense computational power enabled by supercomputers, communication time among processors becomes a bottleneck. Mesh partitioning using METIS [64] ensures optimal load balancing for the computationally intensive simulations. This is crucial for the scalability and efficiency of the model, particularly when simulating complex neurite structures on locally refined truncated T-splines with a large number of DOFs because mesh alterations during execution often lead to unbalanced loads. Unbalanced loads across threads could lead to unnecessary slowdown and waiting during inter-threads/node communications. We ran simulations on the Bridges-2 supercomputer at Pittsburgh Supercomputing Center [103, 104] using 128-thread regular memory nodes.
@@
For clarity, we provide a detailed list of variables for NDDs and healthy neurons in Table 1. The domain $\phi$ is initialized with a central filled circle with radius $r_{0}$, representing the cell. The initial values of $\theta$ are randomized to between $[0,1]$ across the domain with $T$ set to 0 at the beginning. Due to the significant variability in biophysics and growth behaviors across different types of neurons, the parameter values used in the phase field NDDs model were sourced from established literature [29, 30, 105] and fine-tuned empirically to capture realistic neuron growth. These parameters can be adjusted to better simulate biomimetic growth of specific neuron types  [32]. With these parameter settings, we first simulate healthy neuron growth and analyze single- and multiple-neuron growth scenarios (Section 7.1). Then, we conduct a simulation of abnormal neurite morphological transformation using the NDDs model with three case studies that each focuses on a specific parameter (Section 7.2). Finally, we compare simulation results with experimental observations to evaluate our NDDs model (Section 7.3).
@@ 7.1 Healthy neuron growth simulation
In our results, the neuron, represented by $\phi$, is shown in yellow, and the surrounding medium is shown in blue. For single neuron cases, the results (Fig. 4A) show that the model captures a broad spectrum of neurite outgrowth behaviors, from axon elongations to complex branching. Concurrently, our multiple-neuron simulations investigate interactions within complex neurite network formation, varying from 2-7 neurons (Fig. 4B). Starting from 2-neuron cases, we gradually increase complexity to configurations with up to 7 randomly placed neurons. These multi-neuron simulations showcase individual neurite growth patterns and the intricate neurite interactions with responses of other neurons in proximity. With the detailed dynamics of $c_{tubu}$ and $c_{neur}$, these biomimetic single and multi-neuron simulation results showcase the potential of our model for in-depth neurodevelopmental research and provide a possible computational foundation for exploring NDDs.
@@ 7.2 Neurodevelopmental disorders
    Optimal neurotrophin concentration $c_{opti}$. This parameter models the inverse relationship between neurotrophin concentration level and neuron survival and is analyzed for its impact mostly on neurite retraction;
@@
For each study, the parameter being analyzed was adjusted based on the specified ranges listed in Table 1. All other parameters were set according to their default values listed in the same table. By adjusting these parameters in our NDD simulations, we can not only enhance our understanding of the pathophysiological mechanisms of NDDs but also examine the capability of our model to capture the intricate dynamics, which is critical for future application of the NDDs model in developing targeted interventions and advancing neurodevelopmental research.
@@
Optimal neurotrophin concentration $c_{opti}$. By adjusting the value of $c_{opti}$ while keeping other parameters consistent with Table 1, we can better understand its impact on neurite morphological transformations in NDDs. We first analyze neurite growth behaviors under different $c_{opti}$ values ranging from $0.5$ to $3.0$ at around $40,000$ iterations (Fig. 5A). As $c_{opti}$ increases from $0.5$ to $1.5$, there is a noticeable improvement in neurite density and complexity, promoting more extended growth patterns. This process indicates a positive response of neurites to optimal neurotrophin levels, matching our previous expectations of enhanced growth behaviors. As $c_{opti}$ progresses from $1.5$ to $3.0$, there is a noticeable shift towards highly branched and extensively multidirectional neurite structures. This range of growth patterns, from minimal to excessive branching, shows a critical threshold level of neurotrophins necessary for neurodevelopmental processes. Notably, lower values of $c_{opti}$ lead to suppressed neurite outgrowth, validating that there is an inverse relationship between $c_{neur}$ and neuron survival [87], as well as a critical balance of $c_{neur}$ for maintaining healthy neurite development. We then explore the dynamic effects of varying $c_{opti}$ in our NDDs model to understand its impact on neurodevelopmental processes. For simulation cases in Fig. 5B, we initialize $c_{opti}$ as 1 to simulate healthy neurite outgrowth and then reduce $c_{opti}$ to 00 at $150,000$ iterations across cases involving different numbers of neurons. The first two rows in Fig. 5B showcase the growth behavior of individual neurons, showing a straightforward process of neurite retraction as $c_{opti}$ decreases, highlighting the sensitivity of single neurons to $c_{neur}$ and $c_{opti}$ levels. The third and fourth rows extend this analysis to multi-neuron configurations with two and three neurons to investigate neurite retraction and interactions among multiple neurons. These simulation results highlight the intricate influence of $c_{opti}$ levels on the simulated morphology transformation of neurites and the potential implications to retraction and atrophy simulations, vital for informing the future development of effective therapeutic strategies targeting NDDs.
@@
Neurotrophin diffusion rate $D_{c}$. In this study, we focus on the neurotrophin diffusion rate $D_{c}$ that is responsible for controlling the diffusion of $c_{neur}$ to support $\phi$ interface balance. We analyze the effect of $D_{c}$ with varying magnitude from $1.0$ to $6.0$ on single- and multiple-neuron cases, and results at around $100,000$ iterations that best exhibit atrophy are shown in Fig. 6. The results show severe atrophy when the $D_{c}$ value is below $2.0$. As the $D_{c}$ value approaches $3.0$, atrophy becomes moderate. This atrophy behavior indicates that lower values of $D_{c}$, particularly at 1.0, severely restrict the diffusion necessary for generating a neurotrophin gradient field for Eqn. 13 that stabilizes the phase field interface during neurite outgrowth. This insufficient gradient compromises neurite integrity and leads to the degeneration or atrophy of neurite structures due to inadequate neurotrophic support. As a result, as neurite tips continue growing out, there is not enough concentration to support and maintain the grown neurite structures, leading to neurite atrophy. On the other hand, as $D_{c}$ increases, particularly beyond $5.0$, a noticeable improvement in neurite continuity is observed in both single- (Fig. 6A) and multiple-neuron cases (Fig. 6B), markedly reducing the incidence of neurite atrophy. This highlights the necessity of maintaining a diffusion rate $D_{c}$ to ensure a consistent concentration distribution surrounding the simulated growing neurite, thereby supporting the uninterrupted neurodevelopmental process during the simulation and providing insights for potential NDDs progression mitigation.
@@
These detailed parameter studies utilizing our NDDs model demonstrate NDDs morphological transformations to specific simulation parameters. The impact of $c_{opti}$ is most dominantly on neurite retraction and branching, the influence of $D_{c}$ leads to neurite atrophy, and the degradation rates $k_{p75}$ and $k_{2}$ affect the thickness of neurite growth. By categorically investigating these parameters, we have identified crucial factors that predominantly influence the progression of NDDs in our computational model. Our analysis provides crucial insights into how variations in these key parameters can trigger or exacerbate the symptoms associated with NDDs. This knowledge is vital in guiding ongoing research toward understanding the pathophysiological mechanisms underlying these disorders. In the long term, this model could enable more targeted therapeutic planning by focusing on the most influential factors of neuron development and treatment strategy.
7.3 Qualitative validation with experiments
@@
A crucial component of our NDDs model study involves comparing it with experimental observations. We apply the external-cue guided mechanism [32] to simulate the existence of external attractive cues in the extracellular matrix for several simulation cases and set the results against experimental neuron growth patterns (Fig. 8-10). By selectively placing external cues around the neuron, our model can effectively guide neurite outgrowth toward these external cues, allowing it to capture the complex and dynamic growth patterns observed in experimental cultures. These simulations explore how well our model reflects real-world neural growth behaviors. This study compares simulation results with single neuron cultures to better understand the effect of model parameters on individual neurite morphology. Although culturing denser neuron cultures is feasible, the densely overlapping neurite structures will significantly complicate neurite identifications and comparisons. Considering these limitations, efforts have been made to develop sparser neuron cultures to improve the clarity and reliability of neurite analysis.
@@
Figure 8: Comparison of simulated neuron growth patterns and experimental cultures of human iPSC-derived neurons. (A) Comparisons of single neuron cases up to day 5. (B) Comparisons of multiple-neuron cases involving two and three neurons. External cue placements are marked with cyan stars. These comparisons provide a visual qualitative validation of NDDs model to replicate key aspects of neurite morphology and developmental processes seen in vitro.
@@
Comparison with healthy human iPSC-derived neurons. First, we evaluate the model by comparing single-neuron growth patterns from day 1 to day 5 (Fig. 8A). Experimental images of human iPSC-derived neurons cultured over five days are used as the reference. We incorporate the external cue-guided mechanism by placing external cues around the soma to direct neurite outgrowth. These results indicate that the NDDs model can capture the observed neurite growth patterns when external cues are placed according to experimental culture images. Next, we test the NDDs model with multiple neuron configurations (Fig. 8B). This setup involved comparing experimental images of multi-neuron growth with simulations where external cues were selectively placed to guide neurite outgrowth. The simulation outcomes closely aligned with the experimental images, demonstrating that the model can biomimetically capture complex growth behaviors in multi-neuron environments. Although measuring the exact biological conditions and translating them into parameters in the phase field model present inherent challenges, this comparison showcases the effectiveness and potential of our NDDs model in capturing essential aspects of the neurodevelopmental process, highlighting its potential as a powerful tool for studying NDDs. These results evaluate the biomimetic capabilities and applicability of the NDDs model in simulating realistic neurodevelopmental processes, providing valuable insights for targeted therapeutic strategies in the future.
@@
Comparison with damaged and degenerating rat hippocampus neurons. Utilizing the external cue placement strategy [32], we perform a detailed comparative analysis of two neuron growth cases with disorders (Fig. 9). This analysis involves modifying the diffusion coefficient $D_{c}$ and the optimal concentration $c_{opti}$. The results are compared with rat hippocampal neuron cultures exhibiting neurite fragmentation morphologies, characterized by observable severance of neurites followed by disintegration. The NDDs model captures atrophy and retraction by adjusting $D_{c}$, which controls the diffusion of neurotrophin concentration, and $c_{opti}$, which sets the optimal neurotrophin concentration for neurodevelopmental processes. Based on experimental images of neuron growth undergoing damage and cell death (Fig. 9A&E). Red outlines are empirically traced based on both experimental observations and accumulated expertise. External cues are placed to guide healthy neurite patterns (Fig. 9B&F). Subsequently, we decrease the magnitude of $D_{c}$ from $6$ to $1$ to simulate reduced diffusion in neurotrophin concentration (Fig. 9C&G). This simulation reveals that neurites suffer significant atrophy and disconnections, showcasing a stark contrast to simulated healthy growth. In the other comparison, we lowered $c_{opti}$ magnitude from $1$ to 00 (Fig. 9D&H). The results show that neurites not only experience significant atrophy but also demonstrate retraction behaviors, which are influenced by the cascading energy balance at the interface. This study highlights the critical influence of $D_{c}$ and $c_{opti}$ on pathological neuron growth affected by NDDs in our model. These parameter adjustments show great potential to enable the NDDs model to simulate a variety of morphological abnormalities and provide insights into the dynamics of NDDs, aiding in the development of future targeted therapies.
@@
Figure 9: Comparison of experimental neuron cultures undergoing neurite degeneration and simulation results. (A&E) Experimental images of neurons exhibiting neurite breaking and fragmentation are highlighted in red outlines. (B&F) Simulated healthy neurons using the NDDs model. (C&G) Simulated neuron morphology with disorders induced by reducing the diffusion rate $D_{c}$ from 6 to 1. (D&H) Simulated neuron morphology with disorders induced by lowering the optimal concentration $c_{opti}$ from 1 to 0.
Figure 10: Additional comparison of experimental rat hippocampal neurons exhibiting dysfunction and simulated growth patterns. (A) Experimental (top) and model (bottom) neurons that experienced mechanical damage during culture. (B) Experimental (top) and model (bottom) neurons that exhibited neurite breaking and fragmentation from culture condition optimization during the first 3-4 days of culture in vitro. These examples demonstrate the ability of the model to capture a wide array of growth patterns and subsequently qualitatively capture the neurite fragmentation and retraction observed in damaged and degenerating neurons.
@@
In addition to the above two in-depth comparison studies, we apply our NDDs model to simulate a range of cases following the same approach (Fig. 10). The first and third rows in Fig. 10 are the experimental images of neurons exhibiting degeneration and neurite fragmentation (in red outlines), following mechanical damage or culture condition optimization, respectively. The second and fourth rows in Fig. 10 are the corresponding NDDs model simulation results of NDDs neurite growth patterns. Each simulation is able to simulate biomimetic growth behaviors under non-ideal conditions with atrophy and retraction. The ability of the NDDs model to qualitatively capture the neurite fragmentation and retraction observed during cell death and degeneration through biologically plausible parameter tuning highlights its potential for use in modeling NDDs. Future studies should calibrate parameters to individual disorders researchers wish to study.